# Chen, Simchi-Levi, and Wang

### v2

@@ 4.1 Lower Bound for Experimentation with General Treatment
This lower bound effectively implies that asymptotic variance of A/B testing cannot be further reduced under general treatment, which is of independent interest to those who are considering experiments on discounted reward MDPs. In recent literature, Farias et al. (2022) derive a lower bound for A/B testing with general treatment on average-reward MDPs. However, their approach differs from ours in several key aspects. Firstly, their definition of ATE contrasts with ours; they define it as the difference between average rewards for $\pi^{\mathfrak{t}}$ and $\pi^{\mathfrak{c}}$ under stationary distribution, i.e. ${\bm{\mu}}^{{\mathfrak{t}}^{\top}}{\bm{r}}^{\mathfrak{t}}-{\bm{\mu}}^{{\mathfrak{c}}^{\top}}{\bm{r}}^{\mathfrak{c}}$. Additionally, our study further incorporates reward variance, whereas they assume that the reward observation is noise-free. From a technical perspective, our approach employs a distinct proof framework. We apply a constrained Cramér-Rao bound to establish the lower bound of variance for the transition matrix estimation. This represents a departure from the methodology used in Farias et al. (2022).
@@ 5 Variance Reduction: Information Sharing between Test Arms
In this section, we explore how the structure of SST can be leveraged to reduce variance. The underlying principle is that, thanks to the Markov property, the state transitions and reward signals of states $s_{t}\neq s^{1}$ can be shared across different test arms. To illustrate this, let us revisit the trajectory in Section 4. Although $(s^{3},{\mathfrak{c}},s^{3},r_{3})$ is generated by $\pi^{\mathfrak{c}}$, it can also be seen as having been generated by $\pi^{\mathfrak{t}}$ since $\pi^{\mathfrak{t}}(s^{3})=\pi^{\mathfrak{c}}(s^{3})={\mathfrak{c}}$. This observation suggests a potential strategy to use such samples for estimating both ${\bm{V}}^{\mathfrak{t}}$ and ${\bm{V}}^{\mathfrak{t}}$. Consequently, we recolor the trajectory with additional purple on the samples that can be shared between the two policies.
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(s^{1},{\mathfrak{t}},s^{3},r_{1})}\quad{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}(s^{3},{\mathfrak{c}},s^{3},r_{3})}\quad{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}(s^{3},{\mathfrak{c}},s^{1},r_{3})}\quad{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}(s^{1},{\mathfrak{c}},s^{2},r_{4})}\quad{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}(s^{2},{\mathfrak{c}},s^{4},r_{5})}\quad{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}(s^{4},{\mathfrak{c}},s^{5},r_{6})}\cdots$
@@
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\mathcal{D}}^{\mathfrak{t}}_{\textrm{IS}}}=\{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(s_{i},a_{i},s_{i+1},r_{i})}:(s_{i},a_{i})=(s^{1},{\mathfrak{t}})\}\cup\{{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}(s_{i},a_{i},s_{i+1},r_{i})}:s_{i}\not=s^{1}\},$
${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}{\mathcal{D}}^{\mathfrak{c}}_{\textrm{IS}}}=\{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}(s_{i},a_{i},s_{i+1},r_{i})}:(s_{i},a_{i})=(s^{1},{\mathfrak{c}})\}\cup\{{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}(s_{i},a_{i},s_{i+1},r_{i})}:s_{i}\not=s^{1}\}.$
Applying either Algorithm 1 or Algorithm 2 on the datasets ${\mathcal{D}}^{\mathfrak{t}}_{\textrm{IS}}$ and ${\mathcal{D}}^{\mathfrak{c}}_{\textrm{IS}}$ yields two new estimators $\widehat{\bm{V}}^{\mathfrak{t}}_{\textrm{IS}}$ and $\widehat{\bm{V}}^{\mathfrak{c}}_{\textrm{IS}}$ respectively. Subsequently, the estimator for ATE, denoted as $\widehat{\bm{\Delta}}^{\mathrm{IS}}$, is calculated as $\widehat{\bm{V}}^{\mathfrak{t}}_{\textrm{IS}}-\widehat{\bm{V}}^{\mathfrak{c}}_{\textrm{IS}}$. We will always use “IS” to denote information sharing. Observing any given trajectory $\tau$, it is evident that ${\mathcal{D}}^{\mathfrak{t}}_{\textrm{AB}}\subseteq{\mathcal{D}}^{\mathfrak{t}}_{\textrm{IS}}$ and ${\mathcal{D}}^{\mathfrak{c}}_{\textrm{AB}}\subseteq{\mathcal{D}}^{\mathfrak{c}}_{\textrm{IS}}$. This suggests that information sharing is beneficial as it provides a larger dataset for fitting both ${\bm{V}}^{\mathfrak{t}}$ and ${\bm{V}}^{\mathfrak{c}}$ compared to the general A/B testing in Algorithm 3. However, one caveat here is that $\widehat{\bm{V}}^{\mathfrak{t}}_{\textrm{IS}}$ and $\widehat{\bm{V}}^{\mathfrak{c}}_{\textrm{IS}}$ may be sharing a significant portion of the data, introducing correlations between the estimations.
@@
Assuming the Markov chain induced by the experiment policy $\pi^{1/2}$ is uniformly ergodic, with SST, it holds that as $T\to\infty$,
[…]
where $\Sigma_{\mathrm{IS}}$ is provided in Appendix 14.1. Specifically, for the crucial state $s^{1}$, it holds that as $T\to\infty$,
[…]
@@ 5.1 Analysis of Variance Reduction with Two Test Arms
$\sigma_{\mathrm{AB}}^{2}(s^{1})$ $=2\sum_{s\in{\mathcal{S}}}\frac{\rho^{\mathfrak{t}}_{s^{1}}(s)^{2}}{\mu_{1/2}(s)}\Big{(}\sigma_{\mathfrak{t}}^{2}(s)+\left\|{\gamma{\bm{V}}^{\mathfrak{t}}}\right\|_{\Sigma_{p^{\mathfrak{t}}}(s)}^{2}\Big{)}+2\sum_{s\in{\mathcal{S}}}\frac{\rho^{\mathfrak{c}}_{s^{1}}(s)^{2}}{\mu_{1/2}(s)}\left({\sigma_{\mathfrak{c}}^{2}(s)+\left\|{\gamma{\bm{V}}^{\mathfrak{c}}}\right\|_{\Sigma_{p^{\mathfrak{c}}}(s)}^{2}}\right)$
$=\underbrace{2\left({\frac{\rho^{\mathfrak{t}}_{s^{1}}(s^{1})^{2}}{\mu_{1/2}(s^{1})}\Big{(}\sigma_{\mathfrak{t}}^{2}(s^{1})+\left\|{\gamma{\bm{V}}^{\mathfrak{t}}}\right\|_{\Sigma_{p^{\mathfrak{t}}}(s^{1})}^{2}\Big{)}+\frac{\rho^{\mathfrak{c}}_{s^{1}}(s^{1})^{2}}{\mu_{1/2}(s^{1})}\left({\sigma_{\mathfrak{c}}^{2}(s^{1})+\left\|{\gamma{\bm{V}}^{\mathfrak{c}}}\right\|_{\Sigma_{p^{\mathfrak{c}}}(s^{1})}^{2}}\right)}\right)}_{\Sigma^{\mathrm{het}}_{\mathrm{AB}}}$
@@
The $\Sigma^{\mathrm{het}}_{\mathrm{AB}}$ part is the variance directly related to the crucial state $s^{1}$, and the superscript het reflects that state $s^{1}$ is the only state where $\pi^{\mathfrak{t}}$ and $\pi^{\mathfrak{c}}$ may be heterogeneous under SST. The $\Sigma^{\mathrm{hom}}_{\mathrm{AB}}$ part represents the variance associated with all other states $s\neq s^{1}$ and $\pi^{\mathfrak{t}}$ and $\pi^{\mathfrak{c}}$ are homogeneous. We want to remind that under SST, the variances $\sigma^{2}_{\mathfrak{t}}(s)$ and $\sigma^{2}_{\mathfrak{c}}(s)$ are the same for all the other state $s$ except the crucial state $s^{1}$.
@@
where the sum of LHS over $s\neq s^{1}$ constitutes the $\Sigma^{\mathrm{hom}}_{\mathrm{IS}}$ and the sum of RHS forms $\Sigma^{\mathrm{hom}}_{\mathrm{AB}}$. Such a claim leading to the inequality in Eq. (5). To achieve equality in $\Sigma^{\mathrm{hom}}_{\mathrm{IS}}\leq\Sigma^{\mathrm{hom}}_{\mathrm{AB}}$, for all $s\neq s^{1}$, the two Cauthy-Schwarz inequalities above must be equal, which seems to be highly restrictive. In fact, Proposition 5.3 presents the necessary condition for the equality to hold.
@@
Equality in the inequality in Eq. (5) is achieved only if the following two conditions are satisfied simultaneously:
@@
The proof is delayed to Appendix 14.2. This corollary effectively implies that for the equality to hold: either the MDP is trivial with fixed zero reward at every state, or the discounted state visitation measure has to satisfy linear dependence before and after the treatment.
Additionally, we want to highlight that if the new treatment ${\mathfrak{t}}$ is identical to the control ${\mathfrak{c}}$ such that for all states $s$, $\rho^{\mathfrak{t}}_{s^{1}}(s)=\rho^{\mathfrak{c}}_{s^{1}}(s)$ and ${\bm{V}}^{\mathfrak{t}}={\bm{V}}^{\mathfrak{c}}$, the term $\Sigma^{\mathrm{hom}}_{\mathrm{IS}}$ will be zero. In contrast, $\Sigma^{\mathrm{hom}}_{\mathrm{AB}}$ is a value that can be strictly large than zero. This finding is particularly relevant in contexts like A/A testing. Such an observation is somewhat surprising. With information sharing, the amount of data used to estimate the transition probabilities and the rewards can be thought to be doubled, and thus it seems reasonable to conjecture that IS can at most reduce $\Sigma^{\mathrm{hom}}_{\mathrm{AB}}$ to a half, i.e., $\Sigma^{\mathrm{hom}}_{\mathrm{IS}}\geq\frac{1}{2}\Sigma^{\mathrm{hom}}_{\mathrm{AB}}$. However, with the example of A/A testing, the conjecture may not always hold true, suggesting that $\Sigma^{\mathrm{hom}}_{\mathrm{IS}}$ could sometimes be significantly smaller than $\Sigma^{\mathrm{hom}}_{\mathrm{AB}}$.
5.2 Lower Bound for Experimentation with SST
In this subsection, we present the variance lower bound of any unbiased estimator taking the SST structure into account, which reflects the intrinsic difficulty of conducting inference with SST.
@@ 5.2 Lower Bound for Experimentation with SST
Assume the treatment is SST with two test arms. If running mixed policy $\pi^{1/2}$ with initial distribution $\mu_{1/2}$, for any unbiased estimator of ATE ${\bm{\Delta}}$ with covariance matrix $\Sigma$, it holds $\Sigma_{\mathrm{IS}}\preceq T\Sigma$, where $\Sigma_{\mathrm{IS}}$ is defined in Theorem 5.2.
@@
After collecting data, the traditional A/B/n testing would split the trajectory according to different polices and obtain dataset ${\mathcal{D}}^{j}_{\mathrm{ABn}}\coloneqq\{(s_{i},a_{i},s_{i+1},r_{i}):\pi_{i}=\pi^{j},i\in[T]\}$ for each policy $\pi^{j}$. With information sharing, the observations with $s_{i}\neq s^{1}$ among all $n$ test arms can be shared for all SST polices $\pi$. Then, IS yields the corresponding dataset ${\mathcal{D}}^{j}_{\mathrm{ISn}}\coloneqq\{(s_{i},a_{i},s_{i+1},r_{i}):(s_{i},a_{i})\neq(s^{1},{\mathfrak{t}}^{k}),\forall k\neq j,i\in[T]\}$ for each policy $\pi^{j}$. Applying Algorithm 1 or Algorithm 2 to fit ${\bm{V}}^{j}$ with ${\mathcal{D}}^{j}_{\mathrm{ABn}}$ and ${\mathcal{D}}^{j}_{\mathrm{ISn}}$ respectively can produce two estimators $\widehat{\bm{V}}^{\textnormal{ABn}}_{j}$ and $\widehat{\bm{V}}^{\textnormal{ISn}}_{j}$. The estimators for the treatment effect become $\widehat{\bm{\Delta}}_{i,j}^{\textnormal{ABn}}=\widehat{\bm{V}}^{\textnormal{ABn}}_{i}-\widehat{\bm{V}}^{\textnormal{ABn}}_{j}$ and $\widehat{\bm{\Delta}}_{i,j}^{\textnormal{ISn}}=\widehat{\bm{V}}^{\textnormal{ISn}}_{i}-\widehat{\bm{V}}^{\textnormal{ISn}}_{j}$. The following Corollary 5.5 generalizes Theorems 4.1 and 5.1 revealing the asymptotic normality of these two estimators on the crucial state $s^{1}$.
@@
[…]
where
$\sigma_{\mathrm{ABn},(i,j)}^{2}(s^{1})$ $=\underbrace{n\left({\frac{\rho^{i}_{s^{1}}(s^{1})^{2}}{\mu_{1/n}(s^{1})}\left({\sigma_{i}^{2}(s^{1})+\left\|{\gamma{\bm{V}}^{i}}\right\|_{\Sigma_{p^{i}}(s^{1})}^{2}}\right)+\frac{\rho^{j}_{s^{1}}(s^{1})^{2}}{\mu_{1/n}(s^{1})}\left({\sigma_{j}^{2}(s^{1})+\left\|{\gamma{\bm{V}}^{j}}\right\|_{\Sigma_{p^{j}}(s^{1})}^{2}}\right)}\right)}_{\Sigma_{\mathrm{ABn}}^{\mathrm{het}}(i,j)}$
$\qquad+\underbrace{n\sum_{s\neq s^{1}}\frac{\rho^{i}_{s^{1}}(s)^{2}}{\mu_{1/n}(s)}\left({\sigma_{i}^{2}(s)+\left\|{\gamma{\bm{V}}^{i}}\right\|_{\Sigma_{p^{i}}(s)}^{2}}\right)+n\sum_{s\neq s^{1}}\frac{\rho^{j}_{s^{1}}(s)^{2}}{\mu_{1/n}(s)}\left({\sigma_{j}^{2}(s)+\left\|{\gamma{\bm{V}}^{j}}\right\|_{\Sigma_{p^{j}}(s)}^{2}}\right)}_{\Sigma_{\mathrm{ABn}}^{\mathrm{hom}}(i,j)},$
and
$\sigma_{\mathrm{ISn},(i,j)}^{2}(s^{1})=$ $\underbrace{n\left({\frac{\rho^{i}_{s^{1}}(s^{1})^{2}}{\mu_{1/n}(s^{1})}\left({\sigma_{i}^{2}(s^{1})+\left\|{\gamma{\bm{V}}^{i}}\right\|_{\Sigma_{p^{i}}(s^{1})}^{2}}\right)+\frac{\rho^{j}_{s^{1}}(s^{1})^{2}}{\mu_{1/n}(s^{1})}\left({\sigma_{j}^{2}(s^{1})+\left\|{\gamma{\bm{V}}^{j}}\right\|_{\Sigma_{p^{j}}(s^{1})}^{2}}\right)}\right)}_{\Sigma_{\mathrm{ISn}}^{\mathrm{het}}(i,j)}$
$\qquad+\underbrace{\sum_{s\neq s^{1}}\frac{1}{\mu_{1/n}(s)}\left({\big{(}\rho^{i}_{s^{1}}(s)-\rho^{j}_{s^{1}}(s)\big{)}^{2}\sigma_{1}^{2}(s)+\left\|{\gamma\rho^{i}_{s^{1}}(s){\bm{V}}^{i}-\gamma\rho^{j}_{s^{1}}(s){\bm{V}}^{j}}\right\|_{\Sigma_{p^{1}}(s)}^{2}}\right)}_{\Sigma_{\mathrm{ISn}}^{\mathrm{hom}}(i,j)}\,.$

### v3

@@ 3.1 Variance Lower Bound for General A/B Testing on MDPs
This lower bound effectively implies that asymptotic variance of A/B testing cannot be further reduced under general treatment pattern. In recent literature, Farias et al. (2022) derive a similar efficiency result for a similar unbiased A/B testing approach on average-reward MDPs. Their ATE is defined as the difference between average rewards for $\pi^{\mathfrak{t}}$ and $\pi^{\mathfrak{c}}$, i.e. ${\bm{\mu}}^{{\mathfrak{t}}^{\top}}{\bm{r}}^{\mathfrak{t}}-{\bm{\mu}}^{{\mathfrak{c}}^{\top}}{\bm{r}}^{\mathfrak{c}}$. Compared to theirs, our study further incorporates reward variance. From a technical perspective, we employs a distinct proof framework via the constrained Cramér-Rao bound.
@@ 4 Variance Reduction: Information Sharing between Test Arms
In this section, we explore how the structure of SST can be leveraged to reduce variance. The underlying principle is that, thanks to the Markov property, the state transitions and reward signals of states $s_{t}\neq s^{1}$ can be shared across different test arms. To illustrate this, let us revisit the trajectory in Section 3. Although $(s^{3},{\mathfrak{c}},s^{3},r_{2})$ is generated by $\pi^{\mathfrak{c}}$, it can also be seen as having been generated by $\pi^{\mathfrak{t}}$ since $\pi^{\mathfrak{t}}(s^{3})=\pi^{\mathfrak{c}}(s^{3})={\mathfrak{c}}$. This observation suggests a potential strategy to use such samples for estimating both ${\bm{V}}^{\mathfrak{t}}$ and ${\bm{V}}^{\mathfrak{t}}$. Consequently, we recolor the trajectory with additional purple on the samples that can be shared between the two policies.
${\color[rgb]{0,0,1}(s^{1},{\mathfrak{t}},s^{3},r_{1})}\quad{\color[rgb]{.75,0,.25}(s^{3},{\mathfrak{c}},s^{3},r_{2})}\quad{\color[rgb]{.75,0,.25}(s^{3},{\mathfrak{c}},s^{1},r_{3})}\quad{\color[rgb]{1,.5,0}(s^{1},{\mathfrak{c}},s^{2},r_{4})}\quad{\color[rgb]{.75,0,.25}(s^{2},{\mathfrak{c}},s^{4},r_{5})}\quad{\color[rgb]{.75,0,.25}(s^{4},{\mathfrak{c}},s^{5},r_{6})}\cdots$
@@
${\color[rgb]{0,0,1}{\mathcal{D}}^{\mathfrak{t}}_{\textrm{IS}}}=\{{\color[rgb]{0,0,1}(s_{i},a_{i},s_{i+1},r_{i})}:(s_{i},a_{i})=(s^{1},{\mathfrak{t}})\}\cup\{{\color[rgb]{.75,0,.25}(s_{i},a_{i},s_{i+1},r_{i})}:s_{i}\not=s^{1}\},$
${\color[rgb]{1,.5,0}{\mathcal{D}}^{\mathfrak{c}}_{\textrm{IS}}}=\{{\color[rgb]{1,.5,0}(s_{i},a_{i},s_{i+1},r_{i})}:(s_{i},a_{i})=(s^{1},{\mathfrak{c}})\}\cup\{{\color[rgb]{.75,0,.25}(s_{i},a_{i},s_{i+1},r_{i})}:s_{i}\not=s^{1}\}.$
Applying either Algorithm 2 or Algorithm 3 on the datasets ${\mathcal{D}}^{\mathfrak{t}}_{\textrm{IS}}$ and ${\mathcal{D}}^{\mathfrak{c}}_{\textrm{IS}}$ yields two new estimators $\widehat{\bm{V}}^{\mathfrak{t}}_{\textrm{IS}}$ and $\widehat{\bm{V}}^{\mathfrak{c}}_{\textrm{IS}}$ respectively. Subsequently, the estimator for ATE, denoted as $\widehat{\bm{\Delta}}^{\mathrm{IS}}$, is calculated as $\widehat{\bm{V}}^{\mathfrak{t}}_{\textrm{IS}}-\widehat{\bm{V}}^{\mathfrak{c}}_{\textrm{IS}}$. We will always use “IS” to denote information sharing.
@@
Assuming the Markov chain induced by the experiment policy $\pi^{1/2}$ is uniformly ergodic, with SST, it holds that $\sqrt{T}\left({\widehat{\bm{\Delta}}^{\mathrm{IS}}-{\bm{\Delta}}}\right)\xrightarrow{d}{\mathcal{N}}({\bm{0}},\Sigma_{\mathrm{IS}})$ as $T\to\infty$. The detailed form of $\Sigma_{\mathrm{IS}}$ is provided in Appendix 12.4. Specifically, for the crucial state $s^{1}$, it holds that as $T\to\infty$,
$\sqrt{T}\left({\widehat{\Delta}^{\mathrm{IS}}(s^{1})-\Delta(s^{1})}\right)\xrightarrow{d}{\mathcal{N}}(0,\sigma_{\mathrm{IS}}^{2}(s^{1}))\,$,
@@ 4.1 Analysis of Variance Reduction with Two Test Arms
$\sigma_{\mathrm{AB}}^{2}(s^{1})$ $=\underbrace{2\left({\frac{\rho^{\mathfrak{t}}_{s^{1}}(s^{1})^{2}}{\mu_{1/2}(s^{1})}\Big{(}\sigma_{\mathfrak{t}}^{2}(s^{1})+\left\|{\gamma{\bm{V}}^{\mathfrak{t}}}\right\|_{\Sigma_{p^{\mathfrak{t}}}(s^{1})}^{2}\Big{)}+\frac{\rho^{\mathfrak{c}}_{s^{1}}(s^{1})^{2}}{\mu_{1/2}(s^{1})}\left({\sigma_{\mathfrak{c}}^{2}(s^{1})+\left\|{\gamma{\bm{V}}^{\mathfrak{c}}}\right\|_{\Sigma_{p^{\mathfrak{c}}}(s^{1})}^{2}}\right)}\right)}_{\Sigma^{\mathrm{het}}_{\mathrm{AB}}}$
@@
The $\Sigma^{\mathrm{het}}_{\mathrm{AB}}$ part is the variance directly related to the crucial state $s^{1}$, and the superscript het reflects that state $s^{1}$ is the only state where $\pi^{\mathfrak{t}}$ and $\pi^{\mathfrak{c}}$ may be heterogeneous under SST. The $\Sigma^{\mathrm{hom}}_{\mathrm{AB}}$ part represents the variance associated with all other states $s\neq s^{1}$ where $\pi^{\mathfrak{t}}$ and $\pi^{\mathfrak{c}}$ are homogeneous. We want to remind that under SST, the variances $\sigma^{2}_{\mathfrak{t}}(s)$ and $\sigma^{2}_{\mathfrak{c}}(s)$ are the same for all the other state $s$ except the crucial state $s^{1}$.
@@
where the sum of LHS over $s\neq s^{1}$ constitutes the $\Sigma^{\mathrm{hom}}_{\mathrm{IS}}$ and the sum of RHS forms $\Sigma^{\mathrm{hom}}_{\mathrm{AB}}$. Such a claim leading to the inequality in Eq. (4). To achieve equality in $\Sigma^{\mathrm{hom}}_{\mathrm{IS}}\leq\Sigma^{\mathrm{hom}}_{\mathrm{AB}}$, the two Cauthy-Schwarz inequalities above must be equal for all $s\neq s^{1}$, which seems to be highly restrictive. In fact, Proposition 4.3 presents the necessary condition for the equality to hold.
@@
Equality in Eq. (4) is achieved only if the following two conditions are satisfied simultaneously:
@@
The proof is delayed to Appendix 12.4.1. This proposition effectively implies that for the equality to hold: either the MDP is trivial with fixed zero reward at every state, or the discounted state visitation measure has to satisfy linear dependence before and after the treatment.
More importantly, we want to highlight that if the new treatment $\mathfrak{t}$} is identical to the control $\mathfrak{c}$} such that for all states $s$, $\rho^{\mathfrak{t}}_{s^{1}}(s)=\rho^{\mathfrak{c}}_{s^{1}}(s)$ and ${\bm{V}}^{\mathfrak{t}}={\bm{V}}^{\mathfrak{c}}$, the term $\Sigma^{\mathrm{hom}}_{\mathrm{IS}}$ will be zero. In contrast, $\Sigma^{\mathrm{hom}}_{\mathrm{AB}}$ is a value that can be strictly large than zero. This finding is particularly relevant in contexts like A/A testing. Such an observation is somewhat surprising. With information sharing, the amount of data used to estimate the transition probabilities and the rewards can be thought to be doubled, and thus it seems reasonable to conjecture that IS can at most reduce $\Sigma^{\mathrm{hom}}_{\mathrm{AB}}$ to a half, i.e., $\Sigma^{\mathrm{hom}}_{\mathrm{IS}}\geq\frac{1}{2}\Sigma^{\mathrm{hom}}_{\mathrm{AB}}$. However, with the example of A/A testing, the conjecture may not always hold true, suggesting that $\Sigma^{\mathrm{hom}}_{\mathrm{IS}}$ could sometimes be significantly smaller than $\Sigma^{\mathrm{hom}}_{\mathrm{AB}}$.
4.2 Variance Lower Bound for Experimentation with SST
In this subsection, we show that IS, though intuitive and simple, is the optimal way to harness treatment locality for consistent A/B testing. Specifically, we present the variance lower bound of any unbiased estimator taking the SST structure into account, which reflects the intrinsic difficulty of A/B testing with SST.
@@ 4.2 Variance Lower Bound for Experimentation with SST
Assume the treatment is SST with two test arms. If running mixed policy $\pi^{1/2}$ with initial distribution ${\bm{\mu}}_{1/2}$, for any unbiased estimator of ATE $\bm{\Delta}$} with covariance matrix $\Sigma$, it holds $\Sigma_{\mathrm{IS}}\preceq T\Sigma$, where $\Sigma_{\mathrm{IS}}$ is defined in Theorem 4.2.
@@
After collecting data, the traditional A/B/n testing would split the trajectory according to different polices and obtain dataset ${\mathcal{D}}^{j}_{\mathrm{ABn}}\coloneqq\{(s_{i},a_{i},s_{i+1},r_{i}):\pi_{i}=\pi^{j},i\in[T]\}$ for each policy $\pi^{j}$. With information sharing, the observations with $s_{i}\neq s^{1}$ among all $n$ test arms can be shared for all SST polices $\pi$. Then, IS yields the corresponding dataset ${\mathcal{D}}^{j}_{\mathrm{ISn}}\coloneqq\{(s_{i},a_{i},s_{i+1},r_{i}):(s_{i},a_{i})\neq(s^{1},{\mathfrak{t}}^{k}),\forall k\neq j,i\in[T]\}$ for each policy $\pi^{j}$. Applying Algorithm 2 or Algorithm 3 to fit ${\bm{V}}^{j}$ with ${\mathcal{D}}^{j}_{\mathrm{ABn}}$ and ${\mathcal{D}}^{j}_{\mathrm{ISn}}$ respectively can produce two estimators $\widehat{\bm{V}}^{\textnormal{ABn}}_{j}$ and $\widehat{\bm{V}}^{\textnormal{ISn}}_{j}$. The estimators for the treatment effect become $\widehat{\bm{\Delta}}_{i,j}^{\textnormal{ABn}}=\widehat{\bm{V}}^{\textnormal{ABn}}_{i}-\widehat{\bm{V}}^{\textnormal{ABn}}_{j}$ and $\widehat{\bm{\Delta}}_{i,j}^{\textnormal{ISn}}=\widehat{\bm{V}}^{\textnormal{ISn}}_{i}-\widehat{\bm{V}}^{\textnormal{ISn}}_{j}$. The following Corollary 4.5 generalizes Theorems 3.1 and 4.1 revealing the asymptotic normality of these two estimators on the crucial state $s^{1}$.
@@
$\sqrt{T}\left({\widehat{\Delta}_{i,j}^{\Gamma}(s^{1})-\Delta_{i,j}(s^{1})}\right)\xrightarrow{d}{\mathcal{N}}(0,\sigma_{\Gamma,(i,j)}^{2}(s^{1}))\,$,
where $\sigma_{\Gamma,(i,j)}^{2}(s^{1})$ can be decomposed into two terms, the variance associated with the crucial state $\Sigma_{\Gamma}^{\mathrm{het}}(i,j)$ and the one with all the other states $\Sigma_{\Gamma}^{\mathrm{hom}}(i,j)$.